# Aligning LLMs to Ask Good Questions A Case Study in Clinical Reasoning

### v1

@@ 6.3 Preference Tuning Outperforms SFT
6.4 When to integrate the attributes?
We now examine the effect of various attribute integration strategies from §4.3: data-mixing, reward-fusion (PPO only), and policy-fusion. Table 4 reveals three key findings. First, data-mixing achieves higher question win-rate but yields lower diagnostic accuracy than policy-fusion, and reward-fusion in PPO mirrors data-mixing patterns. This suggests that greater attribute separation leads to improved diagnostic accuracy, consistent with prior observations in Yang et al. (2025). Second, The LLM-judge scores strongly correlate with human expert assessments ($r=0.9893,p<0.005$), establishing the LLM-judge as a reliable proxy for human assessment. The high LLM-judge score of Alfa-PPO-Reward is echoed in the human evaluation with a score of 86.3%.
Third, despite high question quality (win-rate), diagnostic accuracy remains suboptimal due to two systemic factors: (i) Patient records derived from r/AskDocs contain only information elicited through human questions. When Alfa-aligned models generate superior questions than that of humans requesting any other useful information, the patient agent lacks corresponding answers. (ii) The final diagnosis generator achieves 86.75% accuracy given complete information, setting an upper bound. This explains why Alfa’s improvements in information-seeking (win-rate) only partially translate to final accuracy. Additionally, performance higher than 86.75% implies that Alfa-aligned models demonstrate emergent reasoning abilities beyond mere data retrieval, producing questions that actively guide clinical reasoning.
*PO POI Size Win-rate MediQ-AD
DPO Data 3B 68.55 85.01
@@ 6.4 When to integrate the attributes?
PPO Data 3B 97.34 84.77
Reward 3B 97.98 84.44
Policy 3B 64.84 86.75
@@
6.5 Gains from Synthetic Data Quality
Model Size Win-rate MediQ-AD
Alfa-Unfiltered 3B 64.19 85.43
8B 63.31 86.09
@@ 6.5 Gains from Synthetic Data Quality
6.6 Synthetic Corruption vs. Enhancement
Figure 4: All the synthetic data directions are helpful. Including corruptions, enhancements, and the original data shows the best performance.
In the counterfactual pairwise data generation stage of Alfa, each original sample $a$ is synthesized in two directions along each attribute dimension (e.g. more relevant and less relevant) to get $a^{+}$ and $a^{-}$. In this section, we examine how the generation direction—enhanced (“more X”) vs. corrupted (“less X”)—influence performance. Specifically, we compare models trained with corruption only pairs $(a,a^{-})$, enhancement only pairs $(a^{+},a)$, synthetic corruption and enhancement pairs $(a^{+},a^{-})$, and all of the above. In Figure 4, we find while any generation direction is beneficial, combining all three pairs brings the most gains, especially apparent in the smaller 3B model.
6.7 Attribute-Specific Influences
@@
Table 6: Policy Fusion DPO models on attribute groups.
Figure 5: Models aligned with general vs. clinical attributes show distinct behaviors. Alfa-General asks clear and focused, but less relevant questions and contains DDX bias ("mono"); Alfa-Clinical uses medical terms hindering answerability but is more professional.
Table 6 shows that removing any attribute leads to performance drops, confirming their importance in clinical question-asking. Clinical attributes have a stronger impact on MediQ accuracy, with avoiding DDX bias being the most critical—suggesting models need explicit training to counteract cognitive biases. Qualitatively, the questions generated by models aligned with general attributes vs. clinical domain-specific attributes show distinct styles (Figure 5), highlighting the impact of feature selection. These results validate Alfa’s structured attribute alignment, demonstrating that both domain-specific capabilities and general question quality contribute to effective clinical decision-making.
6.8 Alfa Models Robustly Generalize to Out-of-Distribution Settings
@@ 7 Related Work
Clinical Reasoning in LLMs.
LLMs have the potential to significantly transform medicine by enhancing personalized care and accessibility (Shanmugam et al., 2024). Models trained with medical data contain rich medical knowledge (Singhal et al., 2025; Lewis et al., 2020; Chen et al., 2023; Labrak et al., 2024; Singhal et al., 2023; Brin et al., 2023), but are limited in instruction-following and multi-hop reasoning Hager et al. (2024); Arroyo et al. (2024); Nov et al. (2023); Zhang et al. (2014). These models excel in static, medical QA benchmarks (Jin et al., 2020; Pal et al., 2022), but recent work has moved away from the static single-turn paradigm and highlight proactive question-asking as a key capability to reliable and effective clinical reasoning (Li et al., 2024; Hu et al., 2024b). Our work furthers this direction by providing an evaluation benchmark with real data and the first method to train specialized question-asking models.
Our proposed method relies on synthetic counterfactual data generation (Mishra et al., 2024a; Ding et al., 2024; Park et al., 2024) and fine-grained alignment. Inspired by prior work on multi-objcetive RLHF (Zhou et al., 2023b; Wu et al., 2023), we extend PPO (Ouyang et al., 2022; Christiano et al., 2017) and DPO (Rafailov et al., 2023) to align models with attribute-specific datasets, and uniquely compare different integration points of the fine-grained preference signals (Rame et al., 2024; Chronopoulou et al., 2023; Wang et al., 2024a).
8 Conclusion
@@
LLM dependence for counterfactual generation. The counterfactual perturbation step relies on LLMs to generate and evaluate counterfactual question variants (specifically, meta-llama/Llama-3.1-405B-Instruct-FP8), assuming they correctly interpret attributes like clarity and relevance. While these attributes are already more concrete than an overall “good question” metric, misinterpretations and evaluation biases remain possible. Future work should incorporate human verification of counterfactuals and attribute-level rankings.
Subjectivity in human annotation. Evaluating follow-up questions is inherently subjective and scenario-dependent. Some annotators expressed difficulty in ranking questions, stating that “none of the questions were good” or that “all of them were acceptable.” Furthermore, we could not require annotators to provide medical certifications, making expertise validation challenging. To mitigate this, we implemented a four-question screening test with known ground-truth answers, filtering out annotators who failed to meet a predefined accuracy threshold. However, this approach does not fully eliminate the risk of variability in domain expertise.
@@
Bias in training data is another key risk. Alfa is trained on online health forum data (r/AskDocs), which may not reflect diverse patient populations, medical conditions, or expert questioning strategies. This could result in systematic biases, where the model generates better questions for certain demographic groups while failing others, reinforcing healthcare disparities. Future work should expand datasets to include diverse patient interactions and implement bias detection and mitigation strategies.
The medical experts recruited as annotators for the task curation and output evaluation are all based in the U.S. and are English-speakers. This not only lead to potential biases in their response, but also assumes US-based medical education background, limiting the generalizability of our results to other countries and regions. The reliance on LLM-judges to score question quality introduces potential automation bias, where models may reinforce subtle inaccuracies or stylistic preferences that diverge from expert human reasoning.
@@
Alfa is intended as a technical contribution to the field of computer science rather than a standalone clinical tool. Even after more comprehensive testing and development, the framework must be integrated with human-in-the-loop supervision, where clinicians retain final decision-making authority. Future work should explore uncertainty calibration, ethical safeguards, and regulatory alignment to ensure safe, fair, and reliable AI-assisted clinical reasoning.
@@ Appendix A Extended Related Works
However, reasoning abilities of these systems, especially under complex, high-stakes demands of medical interaction fulfilling various intentions of users (Nov et al., 2023; Zhang et al., 2014) require further attention. More specifically, clinical reasoning requires ability to ask effective questions (Silverman et al., 2016), crucial for information gathering phase with iterative hypotheses testing and updating. Shaikh et al. (2024) highlighted general lack of question-asking by LLMs in various contexts. While prior works have focused on improving question-asking of LLMs (Andukuri et al., 2024; Rao and Daumé III, 2018) with some focused diagnostic conversations, these works have been limited to rule-based, toy scenarios (Hu et al., 2024b), or prompting-based techniques (Li et al., 2024). Thus, our work expand on such prior works towards a more flexible medical dialogue system with effective question-asking under various real-world user queries.
@@
With the growing capabilities of LLMs and to supplement human data, which can be sparse, synthetic data generation with LLMs have become a popular method (Long et al., 2024; Xu et al., 2024), especially to perform task-specific post-training (Kim et al., 2024). Synthetic data generation is especially appealing in healthcare domain as privacy issues can make data access prohibitive (Ramesh et al., 2024; Xin et al., ; Murtaza et al., 2023). Thus, synthetic data generation through both rule-based methods (Fansi Tchango et al., 2022) and LLMs (Oh et al., 2024; Mishra et al., 2024b; Yao et al., 2024; Wang et al., 2024b) have been explored for various medical tasks. However, data to investigate question quality, especially in medicine, remains under-explored and our data generation method addresses this gap.
@@
To assess LLMs in various medical tasks, many different evaluation frameworks have been developed, typically consisting of static, single-turn question-answering task based on multiple choice questions (Jin et al., 2020; Pal et al., 2022; Jin et al., 2019; Rawat et al., 2024). However, with the advancement of LLMs, there has been a growing need to evaluate LLM agents beyond simple demonstration of knowledge (Thirunavukarasu et al., 2023). MediQ (Li et al., 2024) proposes to evaluate LLMs’ information-seeking ability through interactive clinical reasoning tasks, and constructs a benchmark based on MedQA (Jin et al., 2020) leveraging information asymmetry at benchmark construction time and at inference time. MEDIC (Kanithi et al., 2024) explores comprehensive assessment of LLMs using methods such as LLM-as-a-judge (Zheng et al., 2023). Concurrent work HealthQ (Wang et al., 2024d) analyzes attribute related factors in their evaluation, but lacks theory-grounding and does not propose methods to specifically improve the attributes. Our work builds upon prior works to comprehensively assess LLM’s ability to seek information towards effective clinical communication utilizing LLM-as-a-judge and adopting MediQ framework with a newly generated set of task-specific multiple choice questions (Yao et al., 2025).
@@ Appendix B Dataset Curation
To study clinical conversations, we utilize data from publicly available online health forum r/AskDocs5552013-2021 data accessed at, a subreddit consisting of both lay-users and expert users [^6] Verified by moderators with photo of self with credential documents. See https://www.reddit.com/r/AskDocs/.. We parsed each subsequent comments as a single thread and consider such threads as conversations between users. Since we are interested in clinical followup questions, we first selected threads where the first followup comment from the community contained sentences ending with question marks and further decomposed each conversation into atomic questions, conclusions, and presence of positive feedback from the post author (e.g., thank you) using GPT-4o777gpt-4o-2024-08-06. The resulting dataset contained 17,425 threads, 13,496 unique posts, and 24,263 questions.
@@ C.1 Counterfactual Perturbations
To generate counterfactual perturbations, we use meta-llama/Llama-3.1-405B-Instruct-FP8 with VLLM on 8 A100 80GB GPUs with a temperature of 1.0 and max generation length of 512. See Appendix G.1 for an example prompt.
@@
To verify the quality of generated counterfactual perturbations, we use LLM-judge to rank the generated questions in the perturbed attribute (e.g., accuracy). Furthermore, we used LLM-judge result to filter generated data in § 6.5. To avoid self-preference bias, we used GPT-4o888gpt-4o-2024-08-06. See Appendix G.2 for an example prompt.
In Table 7 below, we report the percentage of pairs kept after the filter in the Enhanced-Corrupted (C-O), Enhanced-Original (E-O), and Original-Corrupted (O-C) directions, as well as the number of training and dev samples after filtering. Intuitively, since the distance between enhanced and corrupted is the largest, the LLM-judge is the most likely to accurately detect the intended perturbation direction, whereas for the Enhanced-Original and Original-Corrupted pairs, more samples fail the LLM-judge filter. We can also see that among all the attributes, clarity has the lowest data quality before filtering.

### v2

@@ 6.3 Preference Tuning Outperforms SFT
     *PO POI Size Win-rate MediQ-AD Win-rate (human)
     DPO Data 3B 68.55 85.01 52.00
@@
     PPO Data 3B 97.34 84.77 75.00
Reward 3B 97.98 84.44 74.00
Policy 3B 64.84 86.75 40.00
@@
Model Model Win MediQ-AD
Variant Size Rate Acc. (%)
Alfa 3B 64.19 85.43
-Unfiltered 8B 63.31 86.09
@@
Figure 4: All synthetic data directions are helpful. Including corruptions, enhancements, and the original data shows the best performance.
6.4 Ablation Studies
I. When to integrate the attributes?
We now examine the effect of various attribute integration strategies from §4.3: data-mixing, reward-fusion (PPO only), and policy-fusion. Table 4 reveals three key findings. First, data-mixing achieves higher question win-rate but yields lower diagnostic accuracy than policy-fusion, and reward-fusion in PPO mirrors data-mixing patterns. This suggests that greater attribute separation leads to improved performance, consistent with prior observations in Yang et al. (2025). Second, The LLM-judge scores had substantial agreement with human expert assessments for strategies with higher win-rate (Gwet’s AC1 score of .68 and .72 each for Alfa-PPO-Reward and Alfa-PPO-Data), establishing the LLM-judge as a reliable proxy for human assessment in this simulated interaction environment. The high LLM-judge score of Alfa-PPO-Reward and Alfa-PPO-Data are echoed in the human evaluation with the highest win-rates of 74% and 75% respectively. See Appendix F and G for further details on annotations and qualitative analysis.
II. Gains from Synthetic Data Quality.
III. Synthetic Corruption vs. Enhancement.
In the counterfactual pairwise data generation stage of Alfa, each original sample $a$ is synthesized in two directions along each attribute dimension (e.g. more relevant and less relevant) to get $a^{+}$ and $a^{-}$. In this section, we examine how the generation direction—enhanced (“more X”) vs. corrupted (“less X”)—influence performance. Specifically, we compare models trained with corruption only pairs $(a,a^{-})$, enhancement only pairs $(a^{+},a)$, synthetic corruption and enhancement pairs $(a^{+},a^{-})$, and all of the above. In Figure 4, we find while all directions are beneficial, combining all three pairs brings the most gains, especially apparent in the smaller 3B model.
@@ I. When to integrate the attributes?
Table 6: Policy Fusion DPO w/ attribute groups.
Figure 5: Models aligned with general vs. clinical attributes show distinct behaviors. Alfa-General: clear and focused, but less relevant and contains DDX bias ("mono"); Alfa-Clinical: professional but uses medical terms hindering answerability.
IV. Attribute-Specific Influences.
Table 6 shows that removing any attribute leads to performance drops, confirming their importance in clinical question-asking. Clinical attributes have a stronger impact on MediQ accuracy. Avoiding DDX bias—factors such as premature closure and availability bias leading to incomplete/incorrect differential diagnoses—is the most critical, suggesting that models need explicit training to counteract cognitive biases. Qualitatively, the questions generated by models aligned with general attributes vs. clinical domain-specific attributes show distinct styles (Figure 5), highlighting the impact of feature selection. These results validate Alfa’s structured attribute alignment, demonstrating that both domain-specific capabilities and general question quality contribute to effective clinical decision-making.
V. Alfa Models Robustly Generalize to Out-of-Distribution Settings.
@@
VI. Alfa Outperforms State-of-the-art General & Medical LLMs
Lastly, we compare Alfa-aligned models to both closed-source general purpose models and medical-specific models by plugging these models into MediQ as the question-generator. Table 7 shows that Alfa shows superior downstream task utility, substantially outperforming even much larger SOTA (general and medical) LLMs.
Model Diagnostic Acc. (%)
Alfa-DPO-8B 88.1
GPT-4o 79.8
Gemini-2 79.8
o3-mini 71.2
Alpacare-7B 74.8
Alpacare-13B 73.5
MedAlpaca-13B 68.5
ClinicalCamel-70B 70.5
Meditron-7B 70.2
Meditron-70B 72.5
PMC-LLaMA-7B 71.9
PMC-LLaMA-13B 69.9
Table 7: Diagnostic Accuracy of Various Models
@@ 7 Related Work
LLMs have the potential to significantly transform medicine by enhancing personalized care and accessibility (Shanmugam et al., 2024). Models trained with medical data contain rich medical knowledge (Singhal et al., 2025; Lewis et al., 2020; Chen et al., 2023; Labrak et al., 2024; Singhal et al., 2023; Brin et al., 2023); however, systematic evaluations reveal persistent weaknesses in instruction-following, multi-hop reasoning, and the nuanced pragmatics that arise in real clinical encounters (Hager et al., 2024; Arroyo et al., 2024; Nov et al., 2023; Zhang et al., 2014). These shortcomings are partly masked by the dominance of static, single-turn medical QA benchmarks, on which current models already achieve near-saturated scores (Jin et al., 2020; Pal et al., 2022). Recent work has moved away from the static single-turn paradigm and highlight proactive information-seeking as a prerequisite to reliable and effective clinical reasoning (Li et al., 2024; Hu et al., 2024b). CoAD (Wang et al., 2023a) is an early exploration toward interactivity, yet it operates in a markedly simplified symbolic environment: the agent selects from a small closed set of symptoms and diseases under dense supervision—impossible to obtain in real-life interactions. Crucially, the setting eliminates the linguistic realization entirely as the “question” is an index to a symptom list. Alfa tackles the more demanding challenge in the open-text regime: teaching an LLM to decide simultaneously what information remains most diagnostically valuable and how to ask for it in a way that is bias-sensitive and pragmatically appropriate.
Methodologically, our approach builds on recent data-centric alignment techniques that create synthetic preference signals for alignment (Li et al., 2025; Mishra et al., 2024a; Ding et al., 2024; Park et al., 2024). Inspired by prior work on multi-objective RLHF (Zhou et al., 2023b; Wu et al., 2023), we extend PPO (Ouyang et al., 2022; Christiano et al., 2017) and DPO (Rafailov et al., 2023) to align models with attribute-specific datasets, and uniquely compare different integration points of the fine-grained preference signals (Rame et al., 2024; Chronopoulou et al., 2023; Wang et al., 2024a).
8 Discussion
@@
Future Directions.
While Alfa demonstrates significant improvements in clinical question-asking within controlled scenarios, several important directions warrant exploration. First, incorporating contextual factors that shape real clinical reasoning—such as care setting constraints (rural vs. urban hospitals), available diagnostic resources, and physician-patient relationship history—could enhance the framework’s real-world applicability. Future work could explore dynamic weighting mechanisms conditioned on these contexts to improve attribute integration. Additionally, integrating multimodal inputs beyond text, such as tone analysis, non-verbal cues, and existing electronic health record data, could better mirror human clinical interactions. The Alfa framework could also benefit from incorporating collaborative decision-making elements, where models learn to ask patients about their own hypotheses or concerns, and to leverage input from healthcare team members. Finally, extending Alfa to other high-stakes domains requiring systematic information-gathering—such as legal discovery, investigative journalism, or financial risk assessment—could validate its generalizability beyond healthcare while revealing domain-specific attribute requirements.
@@
LLM dependence for counterfactual generation. The counterfactual perturbation step relies on LLMs to generate and evaluate counterfactual question variants (specifically, meta-llama/Llama-3.1-405B-Instruct-FP8), assuming they correctly interpret attributes like clarity and relevance. Future work should incorporate human verification of counterfactuals and attribute-level rankings.
Subjectivity in human annotation. Evaluating follow-up questions is inherently subjective and scenario-dependent. Some annotators expressed difficulty in ranking questions, stating that “none of the questions were good” or that “all of them were acceptable.” To ensure annotation quality, we implemented a four-question screening test with known ground-truth answers, filtering out annotators who failed to meet a predefined accuracy threshold. However, this approach does not fully eliminate the risk of variability in domain expertise.
@@
The other presumption underlying this work is that all of these medical problems have verifiable solutions. In actual medical practice—particularly in settings where problems are acute and undefined—it is quite likely that clinicians will come to slightly or entirely different solutions for the same problem (based on their own experiences or expertise).
@@
Bias in training data and evaluation is a key risk. Alfa, trained on r/AskDocs, may not represent diverse populations, conditions, or expert strategies, leading to systematic biases that reinforce healthcare disparities. U.S.-based, English-speaking annotators further limit generalizability. Reliance on LLM-judges may introduce automation bias, reinforcing subtle inaccuracies. Future work should expand datasets and evaluation to more diverse populations and implement bias mitigation strategies.
@@
Alfa is intended as a technical contribution to the field of computer science rather than a standalone clinical tool. The framework must be integrated with human-in-the-loop supervision, where clinicians retain final decision-making authority. Future work should explore uncertainty calibration, ethical safeguards, and regulatory alignment to ensure safe, fair, and reliable AI-assisted clinical reasoning.
Acknowledgements
This research was developed with funding from the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the author and do not reflect the official policy or position of the Department of Defense or the U.S. Government. This material is based upon work supported by the Defense Advanced Research Projects Agency and the Air Force Research Laboratory, contract number(s): FA8650-23-C-7316. Any opinions, findings and conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of AFRL or DARPA. We would like to thank Dr. Saeed Hassanpour and the Dartmouth Center for Precision Health and Artificial Intelligence (CPHAI) for helping to facilitate our collaboration with the medical professionals team in partnership with Lavita AI.
@@ Appendix A Extended Related Works
Reasoning abilities of these systems, especially under complex, high-stakes demands of medical interaction fulfilling various intentions of users (Nov et al., 2023; Zhang et al., 2014) require further attention. More specifically, clinical reasoning requires ability to ask effective questions (Silverman et al., 2016), crucial for information gathering phase with iterative hypotheses testing and updating. Shaikh et al. (2024) highlighted general lack of question-asking by LLMs in various contexts. While prior works have focused on improving question-asking of LLMs (Andukuri et al., 2024; Rao & Daumé III, 2018) with some focused diagnostic conversations, these works have been limited to rule-based, toy scenarios (Hu et al., 2024b), or prompting-based techniques (Li et al., 2024). Further, ALBA (Varadarajan et al., 2024) examined question-asking under mental health assessment setting. Thus, our work expand on such prior works towards a more flexible medical dialogue system with effective question-asking under various real-world user queries.
@@
With the growing capabilities of LLMs and to supplement human data, which can be sparse, synthetic data generation with LLMs have become a popular method (Long et al., 2024; Xu et al., 2024), especially to perform task-specific post-training (Kim et al., 2024). Synthetic data is especially appealing in healthcare domain as privacy issues can make data access prohibitive (Ramesh et al., 2024; Xin et al., ; Murtaza et al., 2023). Thus, synthetic data generation through both rule-based methods (Fansi Tchango et al., 2022) and LLMs (Oh et al., 2024; Mishra et al., 2024b; Yao et al., 2024; Wang et al., 2024b) have been explored for various medical tasks. However, data to investigate question quality, especially in medicine, remains under-explored and our data generation method addresses this gap.
@@
To assess LLMs in various medical tasks, many different evaluation frameworks have been developed, typically consisting of static, single-turn question-answering task based on multiple choice questions (Jin et al., 2020; Pal et al., 2022; Jin et al., 2019; Rawat et al., 2024). However, with the advancement of LLMs, there has been a growing need to evaluate LLM agents beyond simple demonstration of knowledge (Thirunavukarasu et al., 2023). MediQ (Li et al., 2024) proposes to evaluate LLMs’ information-seeking ability through interactive clinical reasoning tasks, and constructs a benchmark based on MedQA (Jin et al., 2020) leveraging information asymmetry at benchmark construction time and at inference time. MEDIC (Kanithi et al., 2024) explores comprehensive assessment of LLMs using methods such as LLM-as-a-judge (Zheng et al., 2023). Concurrent work HealthQ (Wang et al., 2024d) analyzes attribute related factors in their evaluation, but lacks theory-grounding and does not propose methods to specifically improve the attributes. Our work builds on prior works to comprehensively assess LLM’s ability to seek information towards effective clinical communication utilizing LLM-as-a-judge and adopting MediQ framework with a newly generated set of task-specific multiple choice questions (Yao et al., 2025).
@@ Appendix B Dataset Curation
To study clinical conversations, we utilize data from publicly available online health forum r/AskDocs6662013-2021 data accessed at, a subreddit consisting of both lay-users and expert users [^7] Verified by moderators with photo of self with credential documents. See https://www.reddit.com/r/AskDocs/.. We parsed each subsequent comments as a single thread and consider such threads as conversations between users. Since we are interested in clinical followup questions, we first selected threads where the first followup comment from the community contained sentences ending with question marks and further decomposed each conversation into atomic questions, conclusions, and presence of positive feedback from the post author (e.g., thank you) using GPT-4o888gpt-4o-2024-08-06. The resulting dataset contained 17,425 threads, 13,496 unique posts, and 24,263 questions.
@@
Data Quality.
We employed three key strategies to ensure data quality: (1) r/AskDocs has strict anti-misinformation policies and expert verification processes, (2) previous studies have validated the medical quality of responses in this subreddit, and (3) we filtered for samples with positive feedback and clear conclusions to ensure conversation quality.
@@ C.1 Counterfactual Perturbations
To generate counterfactual perturbations, we use meta-llama/Llama-3.1-405B-Instruct-FP8 with VLLM on 8 A100 80GB GPUs with a temperature of 1.0 and max generation length of 512. See Appendix H.1 for an example prompt.
@@
To verify the quality of generated counterfactual perturbations, we use LLM-judge to rank the generated questions in the perturbed attribute (e.g., accuracy). Furthermore, we used LLM-judge result to filter generated data in § 6.4. To avoid self-preference bias, we used GPT-4o999gpt-4o-2024-08-06. See Appendix H.2 for an example prompt.
In Table 8, we report the percentage of pairs kept after the filter in the Enhanced-Corrupted (C-O), Enhanced-Original (E-O), and Original-Corrupted (O-C) directions, as well as the number of training and dev samples after filtering. Intuitively, since the distance between enhanced and corrupted is the largest, the LLM-judge is the most likely to accurately detect the intended perturbation direction, whereas for the Enhanced-Original and Original-Corrupted pairs, more samples fail the LLM-judge filter. We can also see that among all the attributes, clarity has the lowest data quality before filtering.